OpenStudy (anonymous):

9x+4y=18 will give medals for best answer.

OpenStudy (anonymous):

9x+4y=18 -9x -9x subtract on both sides 4y=-9x+18 4/4= 9/4 + 18/4 divide by 4 y=2.25x+4.5

OpenStudy (anonymous):

Thank you yolo_gogo for helping.

OpenStudy (anonymous):

your welcome if you need any help anytime message me

OpenStudy (anonymous):

ok any time

OpenStudy (anonymous):

no prob did u understand what i did in the problem?